Update for the Week Ending October 7, 2016
Site Remediation Continues
- Continued the application of water, odor-suppressant foam, and BioSolve for odor and dust control;
- Maintained erosion and sediment (E&S) controls (Installed waddles - wood chip and straw filled fabric tubes - around the Site perimeter to prevent silt and sediment from leaving the Site and covered all stockpiles);
- Continued pumping from the decontamination pads, excavated areas, sumps and dewatering wells to the water treatment plant (WTP);
- Continued operation of the WTP and discharge to the sewer;
- Continued shallow excavation on the western areas of the Site;
- Continued backfilling and compacting select fill in the shallow excavation in the western areas of the Site;
- Continued lining and loading trucks and shipping impacted soil to Bayshore Soil Management for treatment and disposal:
- 55 trucks; 1,559.89 tons shipped last week; and
- 29,878.46 tons shipped to date.
- Continued the installation of tie-backs through the SOE sheet pile wall for external bracing:
- 11 tie-backs installed last week; and
- 24 of 76 tie-backs installed to date.
- Initiated testing of the installed tie-backs.
- Continue deep excavation within the SOE wall under the TCB.
Community Air Monitoring Plan Summary
Community air monitoring, initiated on Tuesday, December 1, 2015, continues as intrusive (ground disturbance) activities are conducted.
There were no fifteen-minute Action Limit exceedances for or odors during this reporting period.
There was an action level exceedance for Particulate Matter (PM-10) [dust] on October 3, 2016 at AMS-1 during shallow excavation and breaking of subsurface concrete structures while water was being applied to the excavation. Additional water was applied, and PM levels dropped below the response limit.
There was an action level exceedance for Total Volatile Organic Compounds (TVOC) on October 3, 2016 at AMS-2 during fueling operations (diesel and gasoline) for heavy equipment and pressure washers. Once fueling was completed, TVOC levels dropped below the response limit.
There were two action level exceedances for PM-10 on October 6, 2016 at AMS-2 during rock crushing operations. SES applied additional water, and dust levels returned to below the response limit.
If you have any questions or comments about the work on the Site, please use the Williamsburg MGP Telephone Hotline at (718) 403-3053.
Photos of Activity
Installing (drilling) tie-backs (external bracing) for the SOE wall in the deep excavation (south wall).
Installing (drilling) tie-backs (external bracing) for the SOE wall in the deep excavation (east wall).
Load testing installed tie-backs on the western SOE wall.
Loading clean reuse material from the shallow excavation for transfer to the crushing operation on the eastern area of the Site.
Breaking concrete to be crushed for reuse material and applying water for dust control.
Crushing reuse material on the eastern area of the Site and applying water for dust control.
Excavating impacted material in the shallow excavation.
Excavating impacted material and applying water for dust control.
Applying foam to the shallow excavation for VOC and odor control.
Grading select fill in the shallow excavation area.